In numbers 1-4,

(a) find the critical numbers of the given functions

(b) using either the First or Second Derivative Test, classify whether thd function has a relative maximum or minimum at these critical numbers

(c) find the coordinates of the exterum points

(d) find the points of jnflection

(e) find the x and y - intercepts

(f) summarize the points in a table. include additional points necessary for the graph

(g) sketch the polunomial curves in graphic papers


1. f(x)=x³+x²-x+2

2. f(x)=(x4/4)-(x³/3)-2x²+4x+3

3. f(x)=x4-8x²+9

4. f(x)=x³+6x²+9x+3


1
Expert's answer
2021-11-18T06:53:48-0500

1(a) Critical points of "f(x)=x\u00b3+x\u00b2-x+2"

steps

"find\ f'(x)=3x^2+2x-1"

find X

"Critical \ points\ x=-1, x=\frac{1}{3}"



1(b)"f(x)=x\u00b3+x\u00b2-x+2"

"f'(x)=3x^2+2x-1"

"f'(x)\gt 0:x\lt -1 \ or \ x\gt\ \frac{1}{3}"


"f'(x)\lt0:-1\lt x\lt \frac{1}{3}"




1(c) Coordinates of extrema points "of \ f(x)=x³+x²-x+2"

Local maxima at "x=-1"

"coordinates \ = (-1,3)"


local minima at "x=\frac{1}{3}"

"=(\frac{1}{3},\frac{49}{27})"


1(d)Points of inflection

"f"(x)=6x+2\\find \ where \ f"(x)=0 \\ or \\ undefined: \\ x=-\\frac{1}{3}"

"plug \ x=-\frac{1}{3} \ into \ f(x)=x³+x²-x+2"


"=(-\frac{1}{3},\frac{65}{27} )"


1(e)

"for \ x \ intercepts \ replace\ y\ with\ 0\\ for \ y \ intercepts \ replace\ x\ with\ 0"

"\mathrm{X\:Intercepts}:\:\left(-2,\:0\right),\: \\ \mathrm{Y\:Intercepts}:\:\left(0,\:2\right)"


1(f) plot the points "for f(x)=x\u00b3+x\u00b2-x+2"




1(g) Graph "f(x)=x\u00b3+x\u00b2-x+2"







2(a) "critical \ points\ f\left(x\right)=\left(\frac{x^4}{4}\right)-\left(\frac{x^3}{3}\right)-2x²+4x+3"

"find\ f'(x)=4x^3+3x^2-4x+4\\solve\ for\ x"

"x=-2\\x=1\\x=2"


2(b) first Derivative test for maxima and minimum

"f'(x) = 4x^3-3x^2-4x+4"


"f'(x)\gt0: -2\lt x\lt1\ or \ x\gt2"

"f'(x)\lt0: x\lt-2\ or \ 1\lt x\lt2"





2(c) Coordinates of extrema points

"local\ maxima = (1, \frac{59}{12})\\local\ minima=(-2,-\frac{19}{3} )\\local minima =(2,\frac{13}{3})"


2(d) points of inflection of "f\left(x\right)=\left(\frac{x^4}{4}\right)-\left(\frac{x^3}{3}\right)-2x²+4x+3"

"f"(x)=3x^2-2x-4\\find \ where \ f"(x)=0 \\ or \\ undefined: \\ x=\\frac{1-\\sqrt{13}}{3}, x=\\frac{1+\\sqrt{13}}{3}"


"plug \ x=x=\frac{1+\sqrt{13}}{3} \ and \ x=\frac{1-\sqrt{13}}{3}\ into \ f(x)=\frac{x^4}{4}-\frac{x^3}{3}-2x^2+4x+3"

"=\left(\frac{1-\sqrt{13}}{3},\:-\frac{2\left(61+35\sqrt{13}\right)}{81}+3\right),\:\left(\frac{1+\sqrt{13}}{3},\:\frac{2\left(35\sqrt{13}-61\right)}{81}+3\right)"


2(e)

"for \ x \ intercepts \ replace\ y\ with\ 0\\ for \ y \ intercepts \ replace\ x\ with\ 0"

"X \ intercepts=\left(-0.59732\dots ,\:0\right),\:\left(-2.87680\dots ,\:0\right)
\\\:\mathrm{Y\:Intercepts}=\:\left(0,\:3\right)"


2(f) plot the points"f\left(x\right)=\left(\frac{x^4}{4}\right)-\left(\frac{x^3}{3}\right)-2x²+4x+3"




2(g) graph





3(a)Critical points "of \ f(x)=x^4-8x^2+9"

"find\ f'(x)=4x^3-16x\\solve\ for\ x"

"x=-2,\:x=0,\:x=2"


3(b) using first derivative test

"f'(x)= 4x^3-16x\\f'(x)\gt0: -2\lt x\lt0 \ or \ x\gt2\\f'(x)\lt0: x\lt-2 \ or \ 0\lt x\lt2"





3(c) coordinates of extrema points

"local\ maxima = (0, 9)\\local\ minima=(-2,-7 )\\local\ minima =(2,7)"



3(d) points of inflection "of \ f(x)=x^4-8x^2+9"

"f"(x)=12x^2-16\\find \ where \ f"(x)=0 \\ or \\ undefined: \\ x=-\\frac{2\\sqrt{3}}{3}, x=\\frac{2\\sqrt{3}}{3}"


"plug \ x=-2\ into \ f(x)=x^4-8x^2+9=\frac{1}{9}"


"=\quad \left(-\frac{2\sqrt{3}}{3},\:\frac{1}{9}\right),\left(\frac{2\sqrt{3}}{3},\:\frac{1}{9}\right)"


3(e)

"for \ x \ intercepts \ replace\ y\ with\ 0\\ for \ y \ intercepts \ replace\ x\ with\ 0"

"\mathrm{X\:Intercepts}:\:\left(\sqrt{4+\sqrt{7}},\:0\right),\:\left(-\sqrt{4+\sqrt{7}},\:0\right),\:\left(\sqrt{4-\sqrt{7}},\:0\right),\:\left(-\sqrt{4-\sqrt{7}},\:0\right),\\ \:\mathrm{ Y\:Intercepts}:\:\left(0,\:9\right)"



3(f) Plot the points "of \ f(x)=x^4-8x^2+9"





3(g) graph




4(a) critical points of f(x)=x³+6x²+9x+3

"find\ f'(x)=3x^2+6x+9\\solve\ for\ x"


"x=-3,\:x=-1"



4(b) "f'(x)=3x^2+12x^2+9"


"f'(x)\gt0 : x\lt-3\ or\ x\gt-1\\f'(x)\lt0: x\lt x\lt-1"


4(c) coordinates of Extrema points of f(x)=x³+6x²+9x+3

"local\ minima = (-1, -1)\\local\ maxima=(-3,-3)"


4(d) Points of inflection of f(x)=x³+6x²+9x+3

"f"(x)=6x+12\\find \ where \ f"(x)=0 \\ or \\ undefined: \\ x=-2"


"plug \ x=-2\ into \ f(x)=x³+6x²+9x+3=1"


"=\left(-2,\:1\right)"



4(e)

"for \ x \ intercepts \ replace\ y\ with\ 0\\ for \ y \ intercepts \ replace\ x\ with\ 0"

"\mathrm{X\:Intercepts}:\:\left(-0.46791\dots ,\:0\right),\:\left(-1.65270\dots ,\:0\right),\:\left(-3.87938\dots ,\:0\right),\:\\ \mathrm{Y\:Intercepts}:\:\left(0,\:3\right)"



4(f) plot the points of f(x)=x³+6x²+9x+3




4(g) Graph f(x)=x³+6x²+9x+3